# Deep Learning and Machine Learning - Python Data Structures and Mathematics Fundamental: From Theory to Practice

### v1

@@ 1.1 What is Python?
Python is a high-level, interpreted programming language known for its simplicity, readability, and flexibility[138]. It was created by Guido van Rossum and first released in 1991[273]. Python has since become one of the most popular programming languages in the world, used in various domains such as web development, scientific computing, artificial intelligence, and, most importantly for us, mathematical computing. Python is open-source, meaning it’s free to use, and has a large community that contributes to its development and creates powerful libraries for every imaginable use case.
@@ 1.1.1 History and Development of Python
Python’s name comes from “Monty Python’s Flying Circus,”[132] a British sketch comedy show that its creator, Guido van Rossum, enjoyed. He wanted to create a language that was easy and fun to use, while still being powerful enough to solve complex problems. Python has evolved through multiple versions, with Python 2 and Python 3 being the most notable. Python 3 is the current standard, offering many improvements over Python 2.
Guido van Rossum released Python 1.0 in 1994, and since then, Python has continued to grow[272]. Python 3, released in 2008, introduced many changes that were not backward-compatible with Python 2, leading to a gradual shift in adoption by the developer community.
@@ 1.1.3 Setting Up Python: Installation and Environment Configuration
  • PyCharm (for advanced users)[133]
  • VSCode (lightweight and customizable)[61]
  • Jupyter Notebooks (ideal for interactive computing)[139]
@@ 1.3.1 Hello World: The Starting Point
The ”Hello, World!” program is typically the first program written when learning a new programming language[239]. It is very simple and demonstrates the basic syntax of Python.
@@ 2.5 Using Lists, Tuples, Dictionaries, and Sets in Mathematical Computations
Python data structures like lists, tuples, dictionaries, and sets can be very useful for organizing and processing data in mathematical computations[9]. For instance, you can use lists to store numbers, tuples to return multiple values from functions, dictionaries to store key-value pairs of variables and their values, and sets to perform mathematical operations like finding intersections or differences between datasets.
@@ 3.1.1 Control Flow for Decision Making
In Python, we often need to make decisions in the code based on certain conditions[240]. The control flow statements help us define these conditions and control how the code behaves based on the inputs and states of variables. The most commonly used conditional statements in Python are if, elif (short for ”else if”), and else.
@@ 4.1 Introduction to Numpy Arrays
NumPy is a powerful library for numerical computing in Python[92]. It provides support for arrays, matrices, and high-level mathematical functions, which makes it a great choice for performing mathematical operations on large data sets. In this section, we will explore the basics of NumPy arrays and how they differ from Python lists.
@@ 4.2 Matrix Operations in Numpy
In addition to arrays, NumPy also supports matrix operations. Matrices are two-dimensional arrays, and you can perform matrix operations like addition, subtraction, and multiplication[115].
@@ 5.3.1 Dot Product and Cross Product of Vectors
The dot product and cross product of vectors are essential operations in vector algebra[102]. The dot product returns a scalar, while the cross product returns a vector perpendicular to the plane defined by the original vectors.
@@ 5.3.3 Eigenvalues and Eigenvectors
Eigenvalues and eigenvectors are fundamental concepts in linear algebra, particularly useful in systems of equations, stability analysis, and more[204].
@@ 6.1.1 What is Scipy and Why Use It?
Scipy is a Python library used for scientific and technical computing[92]. It builds on the capabilities of NumPy, providing a variety of functions for performing high-level mathematical operations like integration, optimization, interpolation, and many others. Scipy is used for:
@@ 6.1.2 Introduction to Sympy for Symbolic Mathematics
Sympy is a Python library for symbolic mathematics[134]. Unlike Scipy, which focuses on numerical methods, Sympy allows you to perform algebraic manipulations symbolically. This means that you can perform operations like solving algebraic equations, differentiating functions, and integrating expressions exactly, rather than approximately.
@@ 6.3.1 Solving Equations and Limits
This example solves the quadratic equation $x^{2}-4=0$, and solve() returns both solutions: $x=2$ and $x=-2$.
@@ 6.4.1 Understanding the Mathematics of Fourier Transform
The Fourier Transform is a mathematical technique that transforms a time-domain signal into its frequency-domain representation[245]. This is particularly useful for analyzing the frequency content of signals, especially in fields such as signal processing, physics, and engineering.
@@ 6.4.2 Implementing Fourier Transform with Numpy
In Python, we can use the numpy library to perform Fourier Transforms on discrete signals (i.e., sampled data)[276]. The numpy.fft module provides functions to compute the Discrete Fourier Transform (DFT) and its inverse.
@@ 6.4.3 Using FFT for Signal Processing
The Fast Fourier Transform (FFT) is an algorithm that efficiently computes the Discrete Fourier Transform (DFT)[202]. The FFT significantly reduces the computation time compared to the direct calculation of the DFT, making it ideal for real-time signal processing tasks.
@@ 6.5.1 Introduction to Laplace Transform and Applications
The Laplace Transform is a mathematical operation that transforms a function of time $f(t)$ into a function of a complex variable $s$[241]. It is used extensively in the analysis of linear time-invariant (LTI) systems, especially in control systems and circuit analysis.
@@ 6.5.3 Z-Transform in Digital Signal Processing
The Z-Transform is the discrete-time equivalent of the Laplace Transform and is used extensively in digital signal processing (DSP) to analyze and design digital filters and systems[194].
@@ 7.1 Object-Oriented Programming Basics
Object-Oriented Programming (OOP) is a programming paradigm that uses ”objects” to model real-world entities and concepts[287]. Each object is an instance of a class, and classes define the blueprint for these objects. OOP is one of the most powerful ways to write reusable and maintainable code[178].
@@ 8.1.1 Signal Processing: From Time Domain to Frequency Domain
A Fourier Transform breaks down a signal into its constituent frequencies. In Python, this can be achieved using the Fast Fourier Transform (FFT)[203], which is an efficient algorithm for computing the Discrete Fourier Transform (DFT)[252].
@@
  • Sampling Frequency (Fs): This is the number of samples taken per second. In the example, we sample the signal at 500 Hz[277].
@@ 8.1.2 Using FFT to Implement Convolution Efficiently
Convolution is a fundamental operation in signal processing and image processing. It involves combining two signals to form a third signal. Convolution in the time domain can be computationally expensive, especially for large signals. However, the Fourier Transform can be used to compute convolution efficiently[219].
@@
Using the Convolution Theorem[257], we know that convolution in the time domain is equivalent to multiplication in the frequency domain. Here’s how we can implement convolution using FFT:
@@ 8.3 Laplace Transform Applications in Control Systems
The Laplace Transform is an essential mathematical tool used in control systems for analyzing and designing systems in the frequency domain[28]. It transforms a time-domain function into a complex frequency-domain function, making it easier to solve differential equations that describe dynamic systems. In control systems, Laplace transforms help in understanding system behavior, stability, and responses to various inputs.
@@ 8.3.1 Simulating Control Systems with Python
Python, along with libraries like scipy, can be used to simulate control systems and apply the Laplace transform to analyze system responses. One of the key uses of Laplace transforms in control systems is to model and solve linear time-invariant (LTI) systems[80].
@@ 8.4 Comprehensive Project: Frequency Domain Applications in Deep Learning
Deep learning often deals with data in the time or spatial domains, such as audio signals and images[195]. However, transforming this data into the frequency domain using techniques like the Fourier Transform allows for more advanced feature extraction and analysis. This section focuses on how the frequency domain is applied in deep learning tasks, particularly in image processing and audio signal analysis.
@@ 8.4.1 Using Frequency Domain Methods for Image Processing
In image processing, the frequency domain is useful for tasks such as noise reduction, image compression, and feature extraction. The Fourier Transform, specifically the 2D Fourier Transform, is commonly used to analyze and manipulate the frequency components of an image[25, 292].
@@ 8.4.2 Analyzing Audio Signals with Fourier Transform in Deep Learning
In the field of audio signal analysis, the Fourier Transform is used to convert time-domain signals into the frequency domain. This transformation is essential in tasks such as speech recognition, sound classification, and music analysis[22]. Here, we will apply the Fourier Transform to an audio signal and visualize its frequency content.
@@ 9.2.1 Deepening Your Knowledge in Python
2. Practice Coding Challenges: Websites like LeetCode[165], HackerRank[112], and Codewars[56] offer Python challenges that can help you improve your problem-solving skills.
@@ 10.1 Mathematical Foundations in Deep Learning
  • Linear Algebra[109]: Essential for understanding tensors, matrix operations, and vector spaces.
  • Calculus[62]: Necessary for understanding how optimization works (e.g., gradient descent).
  • Probability and Statistics[234]: Important for understanding how models handle uncertainty, interpret data, and measure performance.
  • Optimization[74]: Helps in tuning models to minimize loss functions and improve accuracy.
@@ 10.2 Importance of Linear Algebra and Matrix Operations
Linear algebra plays a pivotal role in deep learning because data is often represented as high-dimensional arrays or matrices (tensors)[55]. Neural networks perform numerous matrix and vector operations such as dot products, matrix multiplication, and element-wise operations, which require a solid understanding of linear algebra.
@@ 10.3 PyTorch and TensorFlow for Mathematical Computations
Both PyTorch and TensorFlow are powerful libraries designed for numerical computation, specializing in handling large-scale tensor operations with automatic differentiation[254]. This makes them ideal for implementing deep learning models.
@@ 11.1 Definition of Tensors
A tensor is a generalization of matrices to higher dimensions. While a scalar is a single number (zero-dimensional), a vector is a one-dimensional array, and a matrix is a two-dimensional array, tensors can be any-dimensional arrays[127]. They are the central data structure in deep learning, representing the data, weights, gradients, and other parameters in a neural network.
@@ 11.4.1 Tensor Initialization (zeros, ones, random)
Initializing tensors is the first step in most deep learning tasks. Common initialization methods include tensors filled with zeros, ones, or random values[196].
@@ 12.1 Element-wise Operations
Element-wise operations are those that are applied to each element of a data structure individually[136]. In Python, we can easily perform element-wise operations on arrays and lists using libraries like NumPy.
@@ 13.2 Optimization of Matrix Multiplication
Matrix multiplication is a fundamental operation in various fields of mathematics and computer science, particularly in linear algebra, computer graphics, machine learning, and numerical analysis[19]. Despite its apparent simplicity, matrix multiplication can be computationally intensive, especially for large matrices. Thus, understanding the intricacies of this operation and exploring optimization techniques is crucial for improving performance in practical applications.
@@ 13.2.2 Traditional Matrix Multiplication
The traditional method of matrix multiplication has a time complexity of $O(n^{3})$, where $n$ is the dimension of the square matrices being multiplied[131]. This cubic complexity arises from the need to compute each entry of the resulting matrix independently, leading to a significant increase in computation time as matrix sizes grow.
@@ 13.2.3 Strassen’s Algorithm
Strassen’s algorithm, introduced by Volker Strassen in 1969, significantly reduces the computational complexity of matrix multiplication from $O(n^{3})$ to approximately $O(n^{2.81})$[261]. This algorithm achieves this reduction by utilizing a divide-and-conquer approach, which minimizes the number of multiplications required.
@@ 13.2.4 Further Improvements in Matrix Multiplication
The Coppersmith-Winograd algorithm further reduced the complexity of matrix multiplication to approximately $O(n^{2.376})$[60]. This algorithm utilizes advanced mathematical techniques involving tensor rank and is considered to be more theoretical due to its complexity and the overhead associated with its practical implementation.
@@
Recent research has led to even faster algorithms, some of which leverage techniques from algebraic geometry and combinatorial optimization. Notably, there have been advancements that utilize fast Fourier transforms (FFT) for multiplying polynomials, which can be adapted to matrix multiplication scenarios, yielding further reductions in complexity[267].
@@ 14.2 LU Decomposition
LU decomposition is a method that decomposes a matrix $A$ into two matrices: a lower triangular matrix $L$ and an upper triangular matrix $U$[108]. This decomposition can simplify the process of solving systems of equations.
@@ 14.3 QR Decomposition
QR decomposition decomposes a matrix $A$ into an orthogonal matrix $Q$ and an upper triangular matrix $R$. This method is particularly useful in solving linear systems and least squares problems[3].
@@ 15.1.2 L2 Norm
The L2 norm (also known as the Euclidean norm) is the square root of the sum of the squares of the vector components[155]. It is defined as:

### v2

@@
   Xinyuan Song
Emory University
xinyuan.song@emory.edu
@@ 1.1 What is Python?
Python is a high-level, interpreted programming language known for its simplicity, readability, and flexibility[jones2020scipy]. It was created by Guido van Rossum and first released in 1991[vanRossum2007Python]. Python has since become one of the most popular programming languages in the world, used in various domains such as web development, scientific computing, artificial intelligence, and, most importantly for us, mathematical computing. Python is open-source, meaning it’s free to use, and has a large community that contributes to its development and creates powerful libraries for every imaginable use case.
@@ 1.1.1 History and Development of Python
Python’s name comes from “Monty Python’s Flying Circus,”[jaworski2019expert] a British sketch comedy show that its creator, Guido van Rossum, enjoyed. He wanted to create a language that was easy and fun to use, while still being powerful enough to solve complex problems. Python has evolved through multiple versions, with Python 2 and Python 3 being the most notable. Python 3 is the current standard, offering many improvements over Python 2.
Guido van Rossum released Python 1.0 in 1994, and since then, Python has continued to grow[vanRossum2003]. Python 3, released in 2008, introduced many changes that were not backward-compatible with Python 2, leading to a gradual shift in adoption by the developer community.
@@ 1.1.3 Setting Up Python: Installation and Environment Configuration
  • PyCharm (for advanced users)[pycharm]
  • VSCode (lightweight and customizable)[vscode]
  • Jupyter Notebooks (ideal for interactive computing)[jupyter]
@@ 1.3.1 Hello World: The Starting Point
The ”Hello, World!” program is typically the first program written when learning a new programming language[sande2019helloworld]. It is very simple and demonstrates the basic syntax of Python.
@@ 2.5 Using Lists, Tuples, Dictionaries, and Sets in Mathematical Computations
Python data structures like lists, tuples, dictionaries, and sets can be very useful for organizing and processing data in mathematical computations[baka2017python]. For instance, you can use lists to store numbers, tuples to return multiple values from functions, dictionaries to store key-value pairs of variables and their values, and sets to perform mathematical operations like finding intersections or differences between datasets.
@@ 3.1.1 Control Flow for Decision Making
In Python, we often need to make decisions in the code based on certain conditions[sanner1999python]. The control flow statements help us define these conditions and control how the code behaves based on the inputs and states of variables. The most commonly used conditional statements in Python are if, elif (short for ”else if”), and else.
@@ 4.1 Introduction to Numpy Arrays
NumPy is a powerful library for numerical computing in Python[fuhrer2021scientific]. It provides support for arrays, matrices, and high-level mathematical functions, which makes it a great choice for performing mathematical operations on large data sets. In this section, we will explore the basics of NumPy arrays and how they differ from Python lists.
@@ 4.2 Matrix Operations in Numpy
In addition to arrays, NumPy also supports matrix operations. Matrices are two-dimensional arrays, and you can perform matrix operations like addition, subtraction, and multiplication[harris2020array].
@@ 5.3.1 Dot Product and Cross Product of Vectors
The dot product and cross product of vectors are essential operations in vector algebra[goldman1985illicit]. The dot product returns a scalar, while the cross product returns a vector perpendicular to the plane defined by the original vectors.
@@ 5.3.3 Eigenvalues and Eigenvectors
Eigenvalues and eigenvectors are fundamental concepts in linear algebra, particularly useful in systems of equations, stability analysis, and more[olver2006applied].
@@ 6.1.1 What is Scipy and Why Use It?
Scipy is a Python library used for scientific and technical computing[fuhrer2021scientific]. It builds on the capabilities of NumPy, providing a variety of functions for performing high-level mathematical operations like integration, optimization, interpolation, and many others. Scipy is used for:
@@ 6.1.2 Introduction to Sympy for Symbolic Mathematics
Sympy is a Python library for symbolic mathematics[johansson2019symbolic]. Unlike Scipy, which focuses on numerical methods, Sympy allows you to perform algebraic manipulations symbolically. This means that you can perform operations like solving algebraic equations, differentiating functions, and integrating expressions exactly, rather than approximately.
@@ 6.3.1 Solving Equations and Limits
This example solves the quadratic equation $x^{2}-4=0$, and solve() returns both solutions: $x=2$ and x=2x=-2.
@@ 6.4.1 Understanding the Mathematics of Fourier Transform
The Fourier Transform is a mathematical technique that transforms a time-domain signal into its frequency-domain representation[sejdic2011fractional]. This is particularly useful for analyzing the frequency content of signals, especially in fields such as signal processing, physics, and engineering.
@@ 6.4.2 Implementing Fourier Transform with Numpy
In Python, we can use the numpy library to perform Fourier Transforms on discrete signals (i.e., sampled data)[Viswanathan2019]. The numpy.fft module provides functions to compute the Discrete Fourier Transform (DFT) and its inverse.
@@ 6.4.3 Using FFT for Signal Processing
The Fast Fourier Transform (FFT) is an algorithm that efficiently computes the Discrete Fourier Transform (DFT)[nussbaumer1982fast]. The FFT significantly reduces the computation time compared to the direct calculation of the DFT, making it ideal for real-time signal processing tasks.
@@ 6.5.1 Introduction to Laplace Transform and Applications
The Laplace Transform is a mathematical operation that transforms a function of time $f(t)$ into a function of a complex variable ss[Schiff2013]. It is used extensively in the analysis of linear time-invariant (LTI) systems, especially in control systems and circuit analysis.
@@ 6.5.3 Z-Transform in Digital Signal Processing
The Z-Transform is the discrete-time equivalent of the Laplace Transform and is used extensively in digital signal processing (DSP) to analyze and design digital filters and systems[nair2004digital].
@@ 7.1 Object-Oriented Programming Basics
Object-Oriented Programming (OOP) is a programming paradigm that uses ”objects” to model real-world entities and concepts[wegner1990concepts]. Each object is an instance of a class, and classes define the blueprint for these objects. OOP is one of the most powerful ways to write reusable and maintainable code[lott2019mastering].
@@ 8.1.1 Signal Processing: From Time Domain to Frequency Domain
A Fourier Transform breaks down a signal into its constituent frequencies. In Python, this can be achieved using the Fast Fourier Transform (FFT)[nussbaumer1982], which is an efficient algorithm for computing the Discrete Fourier Transform (DFT)[smith2007mathematics].
@@
  • Sampling Frequency (Fs): This is the number of samples taken per second. In the example, we sample the signal at 500 Hz[Wahab2016].
@@ 8.1.2 Using FFT to Implement Convolution Efficiently
Convolution is a fundamental operation in signal processing and image processing. It involves combining two signals to form a third signal. Convolution in the time domain can be computationally expensive, especially for large signals. However, the Fourier Transform can be used to compute convolution efficiently[pratt2017fcnn].
@@
Using the Convolution Theorem[stone2008uniqueness], we know that convolution in the time domain is equivalent to multiplication in the frequency domain. Here’s how we can implement convolution using FFT:
@@ 8.3 Laplace Transform Applications in Control Systems
The Laplace Transform is an essential mathematical tool used in control systems for analyzing and designing systems in the frequency domain[broberg1996laplace]. It transforms a time-domain function into a complex frequency-domain function, making it easier to solve differential equations that describe dynamic systems. In control systems, Laplace transforms help in understanding system behavior, stability, and responses to various inputs.
@@ 8.3.1 Simulating Control Systems with Python
Python, along with libraries like scipy, can be used to simulate control systems and apply the Laplace transform to analyze system responses. One of the key uses of Laplace transforms in control systems is to model and solve linear time-invariant (LTI) systems[erfani2013characterisation].
@@ 8.4 Comprehensive Project: Frequency Domain Applications in Deep Learning
Deep learning often deals with data in the time or spatial domains, such as audio signals and images[Najafabadi2015]. However, transforming this data into the frequency domain using techniques like the Fourier Transform allows for more advanced feature extraction and analysis. This section focuses on how the frequency domain is applied in deep learning tasks, particularly in image processing and audio signal analysis.
@@ 8.4.1 Using Frequency Domain Methods for Image Processing
In image processing, the frequency domain is useful for tasks such as noise reduction, image compression, and feature extraction. The Fourier Transform, specifically the 2D Fourier Transform, is commonly used to analyze and manipulate the frequency components of an image[bracewell2012fourier, xu2020learning].
@@ 8.4.2 Analyzing Audio Signals with Fourier Transform in Deep Learning
In the field of audio signal analysis, the Fourier Transform is used to convert time-domain signals into the frequency domain. This transformation is essential in tasks such as speech recognition, sound classification, and music analysis[boashash2015time]. Here, we will apply the Fourier Transform to an audio signal and visualize its frequency content.
@@ 9.2.1 Deepening Your Knowledge in Python
2. Practice Coding Challenges: Websites like LeetCode[leetcode], HackerRank[hackerrank], and Codewars[codewars] offer Python challenges that can help you improve your problem-solving skills.
@@ 10.1 Mathematical Foundations in Deep Learning
  • Linear Algebra[greub2012linear]: Essential for understanding tensors, matrix operations, and vector spaces.
  • Calculus[courant1965introduction]: Necessary for understanding how optimization works (e.g., gradient descent).
  • Probability and Statistics[rohatgi2015introduction]: Important for understanding how models handle uncertainty, interpret data, and measure performance.
  • Optimization[diwekar2020introduction]: Helps in tuning models to minimize loss functions and improve accuracy.
@@ 10.2 Importance of Linear Algebra and Matrix Operations
Linear algebra plays a pivotal role in deep learning because data is often represented as high-dimensional arrays or matrices (tensors)[cichocki2018tensor]. Neural networks perform numerous matrix and vector operations such as dot products, matrix multiplication, and element-wise operations, which require a solid understanding of linear algebra.
@@ 10.3 PyTorch and TensorFlow for Mathematical Computations
Both PyTorch and TensorFlow are powerful libraries designed for numerical computation, specializing in handling large-scale tensor operations with automatic differentiation[stevens2020deep]. This makes them ideal for implementing deep learning models.
@@ 11.1 Definition of Tensors
A tensor is a generalization of matrices to higher dimensions. While a scalar is a single number (zero-dimensional), a vector is a one-dimensional array, and a matrix is a two-dimensional array, tensors can be any-dimensional arrays[hutchison2016lara]. They are the central data structure in deep learning, representing the data, weights, gradients, and other parameters in a neural network.
@@ 11.4.1 Tensor Initialization (zeros, ones, random)
Initializing tensors is the first step in most deep learning tasks. Common initialization methods include tensors filled with zeros, ones, or random values[narkhede2022weight].
@@ 12.1 Element-wise Operations
Element-wise operations are those that are applied to each element of a data structure individually[Johnsson1990]. In Python, we can easily perform element-wise operations on arrays and lists using libraries like NumPy.
@@ 13.2 Optimization of Matrix Multiplication
Matrix multiplication is a fundamental operation in various fields of mathematics and computer science, particularly in linear algebra, computer graphics, machine learning, and numerical analysis[BiniPan2012]. Despite its apparent simplicity, matrix multiplication can be computationally intensive, especially for large matrices. Thus, understanding the intricacies of this operation and exploring optimization techniques is crucial for improving performance in practical applications.
@@ 13.2.2 Traditional Matrix Multiplication
The traditional method of matrix multiplication has a time complexity of $O(n^{3})$, where $n$ is the dimension of the square matrices being multiplied[Iwen2009]. This cubic complexity arises from the need to compute each entry of the resulting matrix independently, leading to a significant increase in computation time as matrix sizes grow.
@@ 13.2.3 Strassen’s Algorithm
Strassen’s algorithm, introduced by Volker Strassen in 1969, significantly reduces the computational complexity of matrix multiplication from $O(n^{3})$ to approximately $O(n^{2.81})$[strassen1969gaussian]. This algorithm achieves this reduction by utilizing a divide-and-conquer approach, which minimizes the number of multiplications required.
@@ 13.2.4 Further Improvements in Matrix Multiplication
The Coppersmith-Winograd algorithm further reduced the complexity of matrix multiplication to approximately $O(n^{2.376})$[coppersmith1987matrix]. This algorithm utilizes advanced mathematical techniques involving tensor rank and is considered to be more theoretical due to its complexity and the overhead associated with its practical implementation.
@@
Recent research has led to even faster algorithms, some of which leverage techniques from algebraic geometry and combinatorial optimization. Notably, there have been advancements that utilize fast Fourier transforms (FFT) for multiplying polynomials, which can be adapted to matrix multiplication scenarios, yielding further reductions in complexity[tolimieri2012mathematics].
@@ 14.2 LU Decomposition
LU decomposition is a method that decomposes a matrix $A$ into two matrices: a lower triangular matrix $L$ and an upper triangular matrix UU[grasedyck2009domain]. This decomposition can simplify the process of solving systems of equations.
@@ 14.3 QR Decomposition
QR decomposition decomposes a matrix $A$ into an orthogonal matrix $Q$ and an upper triangular matrix $R$. This method is particularly useful in solving linear systems and least squares problems[Agarwal2014].
@@ 15.1.2 L2 Norm
The L2 norm (also known as the Euclidean norm) is the square root of the sum of the squares of the vector components[kwak2013principal]. It is defined as: